If a=e^(ix) then (1+a)/(1-a) is equal to
A)cot (x/2) B) tan x C) i cot (x/2) D) i tan x/2 E)2 tan x

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
e^ix = cos x + i sinx 
 
hence 
 
1+a/1-a = 1 + cos x + i sinx / 1 – cos x – i sinx
 
 = (cos x/2 + i sinx/2) / ( sinx/2 – i cosx/2)
 
now you can do it further